You need to make 100 ml of 2% glucose. To do this, you should use 4 ml 50% glucose stock solution and 100 ml water.

True

False
Answer #1

To make 100 ml of 2% glucose from 50% stock solution, we need V1 amount of stock.

C1V1 = C2V2

(50)(V1) = (2)(100)

V1 = 200/50 = 4

We need 4 ml of 50% stock and water to be added will be (100 - 4 = 96 ml).

So, the given statement is false.

Please rate.